OpenStudy (anonymous):

5<3x+8 how do you graph it

OpenStudy (anonymous):

3x - 8 = 5 3x = 5 + 8 3x = 13 (divide both sides by 3 to get x) 3x/3 = 13/3 x = 4.33333333333

OpenStudy (anonymous):

@OpiGeode you did this wrong. it is +8 not -8

OpenStudy (anonymous):

No, look here: https://www.wolframalpha.com/input/?i=5%3C3x%2B8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

she is right

OpenStudy (anonymous):

Who?

OpenStudy (anonymous):

the link even says youre wrong... the correct answer is -1 < x

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Okay.

Nnesha (nnesha):

 this is how to solve first subtract 8 both side then divide by 3 you will end up -1